OpenStudy (anonymous):

Hey, can somebody help me with this problem? :) _____ For the function f(x) = x + 7, what is the ordered pair for the point on the graph when x = 2b? (2b, 2b + 7) (2b, x + 7) (x, x + 7) (x, 2b + 7) _____ Thanks so much guys! <33

jimthompson5910 (jim_thompson5910):

f(x) = x + 7 f(2b) = 2b + 7 ... replace each x with 2b (since x = 2b) and you're done so the ordered pair is (2b, 2b+7)

jimthompson5910 (jim_thompson5910):

remember that y = f(x), so basically the input is x and the output is y or f(x) so plugging in 2b in for x gives you 2b+7 out for y

OpenStudy (anonymous):

Omg you're awesome thank-you so much <33

jimthompson5910 (jim_thompson5910):

you're welcome